Could someone please tell me if my answer is correct?
I have 13.9%
Question is The interest rate "r" to in crease your investemnt "p" to the amount "a" in "t" years is found br "r" = index "t" (radicand a/p)-1. What interest rate would be required to increase your investment of $2700. to $3600. over three years? Thank you

Answer by Alan3354(69443)

The interest rate "r" to in crease your investemnt "p" to the amount "a" in "t" years is found br "r" = index "t" (radicand a/p)-1. What interest rate would be required to increase your investment of $2700. to $3600. over three years?
---------
3600 = 2700*(1 + r)^3 (compounded annually, you didn't specify)
(1+r)^3 = 4/3
1+r =~ 1.10064
--> 10%